OpenStudy (megannicole51):

Find the area inside the cardioid r=1+cos(theta) for 0 is greater than or equal to theta, which is greater than or equal to 2pi.

OpenStudy (megannicole51):

@timo86m .....do you know how to do this? :D

OpenStudy (anonymous):

have you tried this in Polar?

OpenStudy (megannicole51):

what do u mean?

OpenStudy (dumbcow):

here you go: http://en.wikipedia.org/wiki/Polar_coordinate_system#Integral_calculus_.28area.29 \[\rightarrow \frac{1}{2} \int\limits_0^{2\pi}(1+\cos \theta)^{2} d \theta\]

OpenStudy (megannicole51):

((3pi(x))/(2))

OpenStudy (megannicole51):

now what?

OpenStudy (dumbcow):

?? im lost, what is 3pi/2 for

OpenStudy (megannicole51):

value of the integral...

OpenStudy (dumbcow):

oh sorry i was using wolfram and it gave decimal anyway, that is the area

OpenStudy (megannicole51):

my online homework thing just crashed so i cant check the answer....hold please:) lol

OpenStudy (megannicole51):

okay so its 4.71239 :D